You are at a shoe store. There is a pair of shoes for $100. The manager takes 50% off $100.Then he takes 30% off the new price. Finally he takes 20% of the last price. What is the total discount as a percent of the original price. THank You, your help will be appreciated, robert
Answer by stanbon(75887)

Original price = $100
After 50% reduction new price is $50
After 30% reduction new price is 0.7(50)=$35
After 20% reduction new price is 0.8(35)=$28
Total discount is (100-28)/100= 1-0.28=0.72=72%
Cheers,
Stan H.